OpenStudy (anonymous):

given the binomials (x-2),(x-1),(x+2) and (x-4), which one is a factor of f(x)=x^3+7x^2+14x+8

OpenStudy (amistre64):

if its a factor, then they both zero out

OpenStudy (amistre64):

f(x) = (x-r)(other stuff) f(r) = 0

OpenStudy (anonymous):

ok i got (x-4) i wasnt sure if i was right or wrong @amistre64

OpenStudy (amistre64):

if f(4) = 0, its good :) since 4-4 = 0

OpenStudy (amistre64):

4^3+7(4)^2+14(4)+8 got a whole bunch of positives, so i dont think that gets anywhere close to zero; gonna need a negative root at best

OpenStudy (anonymous):

so (x+2)

OpenStudy (amistre64):

yeah, since x+2 is the only negative root; -2+2 = 0, thats the only one id see that would stand a chance

OpenStudy (anonymous):

that would work right?

OpenStudy (anonymous):

ok thnx !

OpenStudy (amistre64):

youre welcome